# Measuring Computational Universality of Fully Homomorphic Encryption

### v1

@@ Abstract.
Many real-world applications, such as machine learning and graph analytics, involve combinations of linear and non-linear operations. As these applications increasingly handle sensitive data, there is a significant demand for privacy-preserving computation techniques capable of efficiently supporting both types of operations—a property we define as "computational universality." Fully Homomorphic Encryption (FHE) has emerged as a powerful approach to perform computations directly on encrypted data. In this paper, we systematically evaluate and measure whether existing FHE methods achieve computational universality or primarily favor either linear or non-linear operations, especially in non-interactive settings. We evaluate FHE universality in three stages. First, we categorize existing FHE methods into ten distinct approaches and analyze their theoretical complexities, selecting the three most promising universal candidates. Next, we perform measurements on representative workloads that combine linear and non-linear operations in various sequences, assessing performance across different bit lengths and with SIMD parallelization enabled or disabled. Finally, we empirically evaluate these candidates on five real-world, privacy-sensitive applications, where each involving arithmetic (linear) and comparison-like (non-linear) operations. Our findings indicate significant overheads in current universal FHE solutions, with efficiency strongly influenced by SIMD parallelism, word-wise versus bit-wise operations, and the trade-off between approximate and exact computations. Additionally, our analysis provides practical guidance to help practitioners select the most suitable universal FHE schemes and algorithms based on specific application requirements. The code is publicly available at https://github.com/UCF-Lou-Lab-PET/FHE-Universality.
Fully Homomorphic Encryption, Universality Measurement
submissionid: 4376copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; ccs: Security and privacy Privacy-preserving protocolsccs: General and reference Measurement\$*\$\$*\$footnotetext: Corresponding Author Email: qian.lou@ucf.edu.
@@ 1. Introduction
FHE is a cryptographic technique that allows computations to be performed directly on encrypted data, eliminating the need for decryption and ensuring sensitive information remains confidential throughout processing. Because of its strong privacy guarantees, FHE has become a cornerstone of secure data analysis in numerous domains, including privacy-preserving machine learning (Xue et al., 2024; Zheng et al., 2023b; Lou and Jiang, 2021), genomics (Kim and Lauter, 2015; Raisaro et al., 2018; Zhang et al., 2015), and finance (Armknecht et al., 2015; Han et al., 2019). It enables researchers and practitioners to extract valuable insights from confidential data while rigorously safeguarding user privacy.
Since Gentry’s first FHE scheme in 2009 (Gentry et al., 2012), numerous constructions (Brakerski, 2012; Brakerski et al., 2014; Cheon et al., 2017; Chillotti et al., 2020; Ducas and Micciancio, 2015; Fan and Vercauteren, 2012; Alexandru et al., 2024; Deng et al., 2024) have been proposed, primarily focusing on enabling either linear or non-linear computations. Existing FHE methods can generally be divided into two categories based on their computational strengths. The first category is word-wise FHE, such as BGV (Brakerski et al., 2014), BFV (Brakerski, 2012; Fan and Vercauteren, 2012), and CKKS (Cheon et al., 2017). These schemes excel at performing efficient word-wise linear operations, such as matrix multiplication, by leveraging Single Instruction Multiple Data (SIMD) (Smart and Vercauteren, 2014) to pack multiple plaintexts into a single ciphertext. The second category is bit-wise FHE, including FHEW (Ducas and Micciancio, 2015) and TFHE (Chillotti et al., 2020), which specialize in non-linear logic operations by supporting arbitrary functions through Boolean circuit evaluations. Bit-wise schemes encrypt each plaintext bit individually, allowing for flexible computation over complex non-linear functions.
However, little attention has been given to systematically enabling mixed linear and non-linear operations in a non-interactive manner. We define this capability as universality. Universality is crucial for many real-world applications that inherently involve a combination of linear and non-linear computations. For example, neural networks (Gilad-Bachrach et al., 2016; Lou et al., 2021) perform linear matrix multiplications followed by non-linear activation functions such as ReLU, while graph algorithms like Floyd-Warshall (Floyd, 1962) often involve linear aggregation and non-linear comparisons.
@@
In addition, given the diverse landscape of FHE methods that are specifically optimized for different applications and workloads, evaluating the FHE’s universality is challenging. Many FHE methods are theoretically capable of universal operations but remain impractical due to significant computational complexity. For instance, directly using bit-wise FHE for universal operations is impractical because of its extremely high computational complexity for linear computation (Lu et al., 2021), i.e., multiplying two 16-bit integers in TFHE can take up to 30 seconds. On the other hand, using word-wise FHE for non-linear functions with linear approximation appears to be a potential direction for handling universal operations, such as polynomial approximation (Cheon et al., 2020, 2019; Lee et al., 2021) for CKKS. However, these methods only work on a small interval, such as $[-1,1]$, inevitably introducing errors around 0. Such errors render approximation-based methods unsuitable for applications where even small errors are intolerable, such as genomics (Kim and Lauter, 2015; Raisaro et al., 2018; Zhang et al., 2015) and finance (Armknecht et al., 2015; Han et al., 2019). Another line of work for universal operations involves scheme switching (Lu et al., 2021; Boura et al., 2020) between word-wise and bit-wise FHE, allowing for the evaluation of the linear operations in word-wise and the non-linear operations in the bit-wise FHE without decryption. While this leverages the strengths of both types, switching cost is prohibitively high, especially for input with large bit-width (Cheon et al., 2022). Most importantly, additional bootstrapping might be needed after scheme switching (Al Badawi et al., 2022), which is one of the most computationally expensive operations in the word-wise FHE and is often avoided in practice (Iliashenko and Zucca, 2021).
@@
Also, many new advancements have emerged and claimed to support non-linear operations using word-wise FHE, but they are not verified in universal computation measurements. For example, word-wise BGV/BFV can perform exact non-linear computations without approximation errors by leveraging polynomial interpolation over the base field $\mathbb{Z}_{p}$ (Tan et al., 2020; Iliashenko and Zucca, 2021; Morimura et al., 2023; Yudha et al., 2024; Zhang et al., 2024b; Xue et al., 2023; Zhang et al., 2023). To enable them to support universal operations in a non-interactive manner, recent work (Zhang et al., 2023) proposed an encoding switching method, HEBridge. Recently, several functional bootstrapping methods (Alexandru et al., 2024; Lee et al., 2024; Liu and Wang, 2023) have been proposed to allow arbitrary function computation during the bootstrapping process. Functional bootstrapping refreshes ciphertexts and applies a chosen function simultaneously, effectively removing noise while completing a specific Look-Up Table (LUT) evaluation. Specifically, Lee et al. (Lee et al., 2024) proposed a functional bootstrapping technique that takes an RLWE ciphertext, i.e., CKKS or BFV, as input and outputs the refreshed BFV ciphertext, building upon BFV-style bootstrapping. However, such BFV-style bootstrapping is far from efficient; for instance, it can take about 3 minutes for a single evaluation due to the intrinsically inefficient slot utilization of BFV-style bootstrapping (Alexandru et al., 2024). To address this issue, Alexandru et al. (Alexandru et al., 2024) presented a general functional bootstrapping technique based on CKKS-style bootstrapping, which has the best throughput among all FHE methods, to improve amortized performance. However, this method suffers from high computational complexity for large input spaces. Their experimental results are limited to 12-bit LUTs because the scaling factor nears the limit for 64-bit modular operations, and the complexity of polynomial evaluation increases significantly, i.e., approximately 1 minute per evaluation for 9-bit LUTs, but about 10 minutes for 12-bit LUTs.
Figure 1. Overview of Our FHE Universality Measurements
@@
To evaluate the computational universality of FHE, we proceed in three stages, as shown in Figure 1. First, we survey and analyze the theoretical complexities of current FHE methods—grouped into 10 distinct directions—to identify three strong candidates for universal computation. Next, we conduct micro-benchmarks on three representative workloads involving mixed linear and non-linear operations: (1) linear followed by non-linear, (2) non-linear followed by linear, and (3) a sequence of linear, non-linear, and linear operations. Finally, we extend our evaluation to five real-world, privacy-sensitive applications that inherently require universal FHE capabilities, including Floyd–Warshall graph analysis, decision tree inference, sorting, database aggregation, and neural network inference. These applications span domains such as finance and healthcare, where protecting sensitive information during complex computations is critical. By integrating theoretical analysis with empirical runtime measurements, we deliver a comprehensive, application-driven benchmark of universal FHE solutions.
@@ 2.1. Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is an encryption technique that enables computations to be carried out directly on encrypted data, ensuring confidentiality throughout processing. Formally, if $y=f(x)$ represents an arithmetic function on plaintext $x$, there exists a function $g(\cdot)$ operating in the encrypted domain such that $y=f(x)\Leftrightarrow y=\mathrm{Dec}(g(\mathrm{Enc}(x)))$. $\mathrm{Enc}(\cdot)$ and $\mathrm{Dec}(\cdot)$ denote encryption and decryption, respectively, while $g(\cdot)$ mirrors $f(\cdot)$ in the ciphertext space. FHE methods introduce noise into ciphertexts to ensure security, typically following the Learning With Errors (LWE) problem. The noise prevents cryptanalysis but must remain below a threshold to ensure successful decryption. The security of modern FHE methods relies on the hardness of LWE and its ring-based variant, Ring LWE (RLWE). In RLWE, the goal is to find the secret $s\in R_{p}$ in the equation $b=a\cdot s+e$, where $b,a\in R_{p}$ are known, and $e$ is an error sampled from a distribution over $R_{p}$. The ring $R_{p}$ is defined as $R_{p}=\mathbb{Z}_{p}[x]/\langle x^{n}+1\rangle$, with $n$ a power of 2 and $p\equiv 1\pmod{2n}$. In practice, ciphertexts are encrypted as tuples of polynomials in this ring, taken modulo an irreducible cyclotomic polynomial whose roots are the $N^{\text{th}}$ primitive roots of unity. The cyclotomic order (degree of the ciphertext polynomials) is typically between $2^{10}$ and $2^{15}$, and the coefficient modulus $q$ is often a product of primes, reaching several hundred bits.
@@
Most current FHE methods can be categorized according to their capability to support different types of operations. Word-wise FHE methods (such as BFV (Fan and Vercauteren, 2012), BGV (Brakerski et al., 2014), and CKKS (Cheon et al., 2017)) excel at linear operations such as matrix multiplication, while bit-wise FHE methods (such as TFHE (Chillotti et al., 2020) and FHEW (Ducas and Micciancio, 2015)) are better suited for non-linear operations such as comparison.
@@
Bit-wise FHE methods. The first category includes bit-wise FHE methods, such as TFHE (Chillotti et al., 2020) and FHEW (Ducas and Micciancio, 2015), which encrypt individual bits into separate ciphertexts, allowing for a wide range of encrypted bitwise manipulations and logic gate evaluations. These schemes support the construction of Boolean circuits composed of encrypted logic gates, facilitating efficient operations directly on encrypted data. TFHE, in particular, is known for its fast bootstrapping, which is essential for evaluating homomorphic logic gates. Bit-wise FHE methods enable developers to leverage decades of digital circuit design research, enabling many applications such as privacy-preserving machine learning and relevant TFHE accelerators (Lou and Jiang, 2019; Lou et al., 2019a; Jiang et al., 2022; Zheng et al., 2023a, 2024; Kumar et al., 2025; Han et al., 2022; Zheng et al., 2022; Lou et al., 2019b). However, they face challenges with addition and multiplication circuits (Lou and Jiang, 2019; Lu et al., 2021), especially when dealing with large circuit depths and fan-in bits. For example, TFHE (Chillotti et al., 2020) takes about 30 seconds to multiply two encrypted 16-bit integers. Also, the expansion ratio of bit-wise FHEs is usually several orders of magnitude larger than that of word-wise FHEs, leading to higher communication costs.
@@
Word-wise FHE methods. Word-wise FHEs can be classified into integer schemes (Brakerski et al., 2014; Brakerski, 2012; Fan and Vercauteren, 2012) and floating-point schemes (Cheon et al., 2017). The two most important FHE schemes that encrypt integers modulo a user-determined modulus $p$ are BGV (Brakerski et al., 2014) and BFV (Fan and Vercauteren, 2012). FHE addition and multiplication correspond to modular addition and multiplication over the plaintext values. It is also possible to emulate Boolean circuits in integer schemes by setting $p=2$, which causes addition to behave as an XOR gate and multiplication to work as an AND gate. Unlike bit-wise FHE schemes, both BGV and BFV have considerably slower bootstrapping; depending on parameter choices (such as the ring dimension), a single bootstrapping can take anywhere from several seconds to several hours (Geelen and Vercauteren, 2023). Thus, most implementations of BGV and BFV do not include bootstrapping and are used exclusively in LHE mode (Gouert et al., 2023).
@@
For the floating-point schemes, they use a plaintext type of floating-point numbers, and the most popular scheme in this category is CKKS (Cheon et al., 2017). It is desirable for certain classes of algorithms, such as machine learning (Xue et al., 2024; Zhang et al., 2025; Zheng et al., 2023c). In practice, the core operations of floating-point schemes parallel those of integer schemes, with only a few differences. Similarly, the floating-point schemes have slow bootstrapping procedures and are predominantly used in LHE mode. The key difference is the need to keep track of the scale factor that is multiplied by plaintext values during encoding, which determines the bit-precision associated with each ciphertext. The scale doubles when two ciphertexts are multiplied, which may result in overflow as the scale becomes exponentially larger. Thus, rescaling must be invoked to preserve the original scale after multiplications, which is similar to modulus switching and serves a similar role in reducing ciphertext noise.
@@
The key advantage of word-wise FHE over bit-wise methods is its support for batching. Single Instruction Multiple Data (SIMD) enables encrypting a vector of plaintexts into a single ciphertext (Smart and Vercauteren, 2014; Cheon et al., 2017). Each plaintext occupies a slot, with the number of slots determined by encryption parameters. Batched ciphertexts support slot-wise addition and multiplication. A key limitation is slot dependency—if operations require interaction across slots (e.g., summing all slots), slot-wise rotations are needed, incurring significant memory and runtime overhead.
@@
Despite their efficiency in linear operations, word-wise FHE methods struggle with non-linear functions like comparison (Lou et al., 2020a, b), prompting efforts to support such functions within word-wise schemes (Iliashenko and Zucca, 2021; Tan et al., 2020). CKKS supports approximate computation on floating-point numbers, and the most prevailing method to perform the non-linear functions is through polynomial approximation (Cheon et al., 2020, 2019; Lee et al., 2021). These methods use a composition of minimax approximation polynomials (Lee et al., 2021) to approximate the sign function with errors. These methods primarily work on only small intervals, such as $[-1,-\epsilon]\cup[\epsilon,1]$ with errors. The smaller the errors are, the higher the degree of polynomial required. Thus, more multiplication and multiplicative depth are needed. Most importantly, the approximation can never be accurate around 00. Such errors limit the application of approximation-based non-linear where accuracy is important, such as genomics (Kim and Lauter, 2015; Raisaro et al., 2018; Zhang et al., 2015) and finance (Armknecht et al., 2015; Han et al., 2019). On the other hand, since computations are exact in BFV and BGV, the non-linear functions can be implemented in BFV and BGV without approximation error. There has been a series of works (Tan et al., 2020; Iliashenko and Zucca, 2021; Morimura et al., 2023) exploring the efficient comparison in BFV and BGV. On a high level, they compute the non-linear functions by evaluating an interpolation polynomial over the base field. However, the degree of the interpolation polynomial grows exponentially with the bit-width of the input (Tan et al., 2020) or relies on a special plaintext space (Iliashenko and Zucca, 2021) that does not support seamless computation of the linear and non-linear operations.
2.2. Universality Required Privacy-sensitive Applications

### v2

# SoK: Can Fully Homomorphic Encryption Support General AI Computation? A Functional and Cost Analysis

@@ Abstract.
Artificial intelligence (AI) increasingly powers sensitive applications in domains such as healthcare and finance, relying on both linear operations (e.g., matrix multiplications in large language models) and non-linear operations (e.g., sorting in retrieval-augmented generation). Fully homomorphic encryption (FHE) has emerged as a promising tool for privacy-preserving computation, but it remains unclear whether existing methods can support the full spectrum of AI workloads that combine these operations.
In this SoK, we ask: Can FHE support general AI computation? We provide both a functional analysis and a cost analysis. First, we categorize ten distinct FHE approaches and evaluate their ability to support general computation. We then identify three promising candidates and benchmark workloads that mix linear and non-linear operations across different bit lengths and SIMD parallelization settings. Finally, we evaluate five real-world, privacy-sensitive AI applications that instantiate these workloads. Our results quantify the costs of achieving general computation in FHE and offer practical guidance on selecting FHE methods that best fit specific AI application requirements. Our codes are available at https://github.com/UCF-ML-Research/FHE-AI-Generality.
Fully Homomorphic Encryption, Generality Measurement
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
The rapid advancement of AI has enabled its widespread adoption across privacy-sensitive domains, including healthcare (kim2015private; raisaro2018protecting; zhang2015foresee) and finance (armknecht2015guide; han2019logistic). To protect privacy, FHE is a promising tool, as it enables direct computation on encrypted data without decryption, ensuring end-to-end confidentiality. However, AI workloads rely on both linear and non-linear operations, while existing FHE schemes are specialized rather than general: word-wise schemes (e.g., BGV (brakerski2014leveled; yudha2024boostcom), BFV (brakerski2012fully; fan2012somewhat), CKKS (cheon2017homomorphic)) are efficient for linear operations but struggle with non-linear ones, whereas bit-wise schemes (e.g., FHEW (ducas2015fhew), TFHE (chillotti2020tfhe)) handle non-linear operations well but incur prohibitively high costs on linear ones. Thus, it remains unclear whether existing FHE can truly support the full spectrum of AI workloads that mix linear and non-linear operations in practice. For example, neural networks (gilad2016cryptonets; lou2021safenet) alternate linear matrix multiplications with non-linear activation functions, and graph algorithms such as Floyd–Warshall (floyd1962algorithm) combine linear aggregation with non-linear comparisons.
We refer to such mixed computations as general computation. The ability of an FHE method to support general computation in a non-interactive and efficient manner is what we call computational generality. Although prior research has systematized FHE along dimensions such as performance, toolchain maturity, and developer accessibility (viand2021sok; gouert2023sok), computational generality has received little attention—despite being essential for real AI applications. The open research question is twofold: first, the functionality of different FHE methods in supporting general computation, and second, the cost of executing it when such support is available.
Figure 1. Overview of Our FHE Generality Measurements.
\$*\$\$*\$footnotetext: Corresponding Author Email: qian.lou@ucf.edu.
@@
Given the diverse landscape of FHE methods that are specifically optimized for different applications and workloads, evaluating the FHE’s computational generality is challenging. Many FHE methods are theoretically capable of general computation but remain impractical due to significant computational complexity. For instance, directly using bit-wise FHE for general computation is impractical because of its extremely high computational complexity for linear computation (lu2021pegasus), i.e., multiplying two 16-bit integers in TFHE can take up to 30 seconds. On the other hand, using word-wise FHE for non-linear functions with linear approximation appears to be a potential direction for handling general computation, such as polynomial approximation (cheon2020efficient; cheon2019numerical; lee2021minimax) for CKKS. However, these methods only work on a small interval, such as [1,1][-1,1], inevitably introducing errors around 0. Such errors render approximation-based methods unsuitable for applications where even small errors are intolerable, such as genomics (kim2015private; raisaro2018protecting; zhang2015foresee) and finance (armknecht2015guide; han2019logistic). Another line of work for general computation involves scheme switching (lu2021pegasus; boura2020chimera) between word-wise and bit-wise FHE, allowing for the evaluation of the linear operations in word-wise and the non-linear operations in the bit-wise FHE without decryption. While this leverages the strengths of both types, switching cost is prohibitively high, especially for input with a large bit-width (cheon2022efficient). Most importantly, additional bootstrapping might be needed after scheme switching (al2022openfhe), which is one of the most computationally expensive operations in the word-wise FHE and is often avoided in practice (iliashenko2021faster).
@@
Additionally, while word-wise BGV/BFV can leverage $\mathbb{F}_{p}$-based polynomial interpolation (tan2020efficient; iliashenko2021faster; morimura2023accelerating) to enable exact non-linear operations, their efficacy in general computations remains unverified, because the inherent field disparity between $\mathbb{F}_{p}$ (non-linear) and $\mathbb{Z}_{p^{d}}$ (linear) fundamentally limits seamless integration of mixed operations. To enable them to support general computation in a non-interactive manner, Zhang et al. (zhang2023hebridge) proposed an encoding switching method, HEBridge. Recently, several functional bootstrapping methods (alexandru2024general; lee2024functional; liu2023amortized) have been proposed to allow arbitrary function computation during the bootstrapping process. Functional bootstrapping refreshes ciphertexts and applies a chosen function simultaneously, effectively removing noise while completing a specific Look-Up Table (LUT) evaluation. Specifically, Lee et al. (lee2024functional) proposed a functional bootstrapping technique that takes an RLWE ciphertext, i.e., CKKS or BFV, as input and outputs the refreshed BFV ciphertext, building upon BFV-style bootstrapping. However, such BFV-style bootstrapping is far from efficient; for instance, it can take about 3 minutes for a single evaluation due to the intrinsically inefficient slot utilization of BFV-style bootstrapping (alexandru2024general). To address this issue, Alexandru et al. (alexandru2024general) presented a general functional bootstrapping technique based on CKKS-style bootstrapping, which has the best throughput among all FHE methods, to improve amortized performance. However, this method suffers from high computational complexity for large input spaces. Their experimental results are limited to 12-bit LUTs because the scaling factor nears the limit for 64-bit modular operations, and the complexity of polynomial evaluation increases significantly, i.e., approximately 1 minute per evaluation for 9-bit LUTs, but about 10 minutes for 12-bit LUTs.
@@
To evaluate the computational generality of FHE, we proceed in three stages, as shown in Figure 1. First, we survey and analyze current FHE methods to determine whether they have the functionality to support general computation. For FHE methods that support general computation, we analyze their theoretical complexities and identify the three most efficient candidates. Next, we conduct micro-benchmarks on three representative general workloads involving mixed linear and non-linear operations: (1) linear followed by non-linear, (2) non-linear followed by linear, and (3) a sequence of linear, non-linear, and linear operations. Finally, we extend our evaluation to five real-world, privacy-sensitive applications that inherently require FHE computational generality, including Floyd–Warshall graph analysis, decision tree inference, sorting, database aggregation, and neural network inference. These applications span domains such as finance and healthcare, where protecting sensitive information during complex computations is critical. By integrating theoretical analysis with empirical runtime measurements, we deliver a comprehensive, application-driven benchmark of general FHE solutions.
@@ 2.1. Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is an encryption technique that enables computations to be carried out directly on encrypted data, ensuring confidentiality throughout processing. Formally, if $y=f(x)$ represents an arithmetic function on plaintext $x$, there exists a function $g(\cdot)$ operating in the encrypted domain such that $y=f(x)\Leftrightarrow y=\mathrm{Dec}(g(\mathrm{Enc}(x)))$. $\mathrm{Enc}(\cdot)$ and $\mathrm{Dec}(\cdot)$ denote encryption and decryption, respectively, while $g(\cdot)$ mirrors $f(\cdot)$ in the ciphertext space. FHE methods introduce noise into ciphertexts to ensure security, typically following the Learning With Errors (LWE) problem. The noise prevents cryptanalysis but must remain below a threshold to ensure successful decryption. The security of modern FHE methods relies on the hardness of LWE and its ring-based variant, Ring LWE (RLWE). In RLWE, the goal is to find the secret $s\in R_{p}$ in the equation $b=a\cdot s+e$, where $b,a\in R_{p}$ are known, and $e$ is an error sampled from a distribution over $R_{p}$. The ring $R_{p}$ is defined as $R_{p}=\mathbb{F}_{p}[x]/\langle x^{n}+1\rangle$, with $n$ a power of 2 and $p\equiv 1\pmod{2n}$. In practice, ciphertexts are encrypted as tuples of polynomials in this ring, taken modulo an irreducible cyclotomic polynomial whose roots are the $N^{\text{th}}$ primitive roots of unity. The cyclotomic order (degree of the ciphertext polynomials) is typically between $2^{10}$ and $2^{15}$, and the coefficient modulus $q$ is often a product of primes, reaching several hundred bits.
@@
Most current FHE methods can be categorized according to their capability to support different types of operations. Word-wise FHE methods (such as BFV (fan2012somewhat), BGV (brakerski2014leveled), and CKKS (cheon2017homomorphic)) excel at linear operations such as matrix multiplication, while bit-wise FHE methods (such as TFHE (chillotti2020tfhe) and FHEW (ducas2015fhew)) are better suited for non-linear operations such as comparison.
@@
Bit-wise FHE methods. The first category includes bit-wise FHE methods, such as TFHE (chillotti2020tfhe) and FHEW (ducas2015fhew), which encrypt individual bits into separate ciphertexts, allowing for a wide range of encrypted bitwise manipulations and logic gate evaluations. These schemes support the construction of Boolean circuits composed of encrypted logic gates, facilitating efficient operations directly on encrypted data. TFHE, in particular, is known for its fast bootstrapping, which is essential for evaluating homomorphic logic gates. Bit-wise FHE methods enable developers to leverage decades of digital circuit design research, enabling many applications such as privacy-preserving machine learning and relevant TFHE accelerators (lou2019she; lou2019glyph; jiang2022matcha; zheng2023priml; zheng2024ofhe; kumar2025tfhe; han2022coxhe; zheng2022cryptolight; zhu2025dahe; lou2019autoq; deng2024trinity). However, they face challenges with addition and multiplication circuits (lou2019she; lu2021pegasus), especially when dealing with large circuit depths and fan-in bits. For example, TFHE (chillotti2020tfhe) takes about 30 seconds to multiply two encrypted 16-bit integers. Also, the expansion ratio of bit-wise FHEs is usually several orders of magnitude larger than that of word-wise FHEs, leading to higher communication costs.
@@
Word-wise FHE methods. Word-wise FHEs can be classified into integer schemes (brakerski2014leveled; brakerski2012fully; fan2012somewhat) and floating-point schemes (cheon2017homomorphic). The two most important FHE schemes that encrypt integers modulo a user-determined modulus $p$ are BGV (brakerski2014leveled) and BFV (fan2012somewhat). FHE addition and multiplication correspond to modular addition and multiplication over the plaintext values. It is also possible to emulate Boolean circuits in integer schemes by setting $p=2$, which causes addition to behave as an XOR gate and multiplication to work as an AND gate. Unlike bit-wise FHE schemes, both BGV and BFV have considerably slower bootstrapping; depending on parameter choices (such as the ring dimension), a single bootstrapping can take anywhere from several seconds to several hours (geelen2023bootstrapping). Thus, most implementations of BGV and BFV do not include bootstrapping and are used exclusively in LHE mode (gouert2023sok).
@@
For the floating-point schemes, they use a plaintext type of floating-point numbers, and the most popular scheme in this category is CKKS (cheon2017homomorphic). It is desirable for certain classes of algorithms, such as machine learning (xue2024cryptotrain; zhang2025cipherprune; zheng2023primer; zhao2025design; zhang2024heprune). In practice, the core operations of floating-point schemes parallel those of integer schemes, with only a few differences. Similarly, the floating-point schemes have slow bootstrapping procedures and are predominantly used in LHE mode. The key difference is the need to keep track of the scale factor that is multiplied by plaintext values during encoding, which determines the bit-precision associated with each ciphertext. The scale doubles when two ciphertexts are multiplied, which may result in overflow as the scale becomes exponentially larger. Thus, rescaling must be invoked to preserve the original scale after multiplications, which is similar to modulus switching and serves a similar role in reducing ciphertext noise.
@@
The key advantage of word-wise FHE over bit-wise methods is its support for batching. Single Instruction Multiple Data (SIMD) enables encrypting a vector of plaintexts into a single ciphertext (smart2014fully; cheon2017homomorphic). Each plaintext occupies a slot, with the number of slots determined by encryption parameters. Batched ciphertexts support slot-wise addition and multiplication. A key limitation is slot dependency—if operations require interaction across slots (e.g., summing all slots), slot-wise rotations are needed, incurring significant memory and runtime overhead.
@@
Despite their efficiency in linear operations, word-wise FHE methods struggle with non-linear functions like comparison (lou2020falcon; lou2020autoprivacy), prompting efforts to support such functions within word-wise schemes (iliashenko2021faster; tan2020efficient). CKKS supports approximate computation on floating-point numbers, and the most prevailing method to perform the non-linear functions is through polynomial approximation (cheon2020efficient; cheon2019numerical; lee2021minimax). These methods use a composition of minimax approximation polynomials (lee2021minimax) to approximate the sign function with errors. These methods primarily work on only small intervals, such as $1,-\epsilon]\cup[\epsilon,1]$ with errors. The smaller the errors are, the higher the degree of the polynomial required. Thus, more multiplication and multiplicative depth are needed. Most importantly, the approximation can never be accurate around 0. Such errors limit the application of approximation-based non-linear methods where accuracy is important, such as genomics (kim2015private; raisaro2018protecting; zhang2015foresee) and finance (armknecht2015guide; han2019logistic). On the other hand, since computations are exact in BFV and BGV, the non-linear functions can be implemented in BFV and BGV without approximation error. There has been a series of works (tan2020efficient; iliashenko2021faster; morimura2023accelerating) exploring the efficient comparison in BFV and BGV. On a high level, they compute the non-linear functions by evaluating an interpolation polynomial over the base field. However, the degree of the interpolation polynomial grows exponentially with the bit-width of the input (tan2020efficient) or relies on a special plaintext space (iliashenko2021faster) that does not support seamless computation of the linear and non-linear operations.
2.2. Bootstrapping Overview
Bootstrapping is a fundamental operation in HE that refreshes noisy ciphertexts to enable unbounded computation. Each homomorphic operation introduces noise into ciphertexts, and once noise exceeds the decryption threshold, correctness fails. Bootstrapping homomorphically evaluates the decryption function on a noisy ciphertext, producing a refreshed ciphertext with reduced noise: $\textbf{ct}^{\prime}=\text{Enc}_{\text{sk}}^{\prime}(\text{Dec}_{\text{sk}}(\textbf{ct}))=\text{Enc}_{\text{sk}}^{\prime}(m)$. The three computationally general FHE methods employ fundamentally different bootstrapping strategies with dramatically different costs. Bit-wise TFHE uses gate bootstrapping (around 102010-20 ms per gate), where each gate evaluation is immediately followed by noise refresh, enabling constant multiplicative depth but requiring separate bootstrapping for each bit. Word-wise methods (Scheme Switching and Encoding Switching) rely on BGV/BFV bootstrapping, which is orders of magnitude slower (3-30 seconds per operation) but can amortize costs across thousands of SIMD slots, achieving  1-7 ms per slot when batching is possible. Due to these prohibitive word-wise bootstrapping costs, most implementations operate in leveled mode with sufficiently large parameters to avoid bootstrapping entirely.
2.3. Generality Required Privacy-sensitive Applications